# Navigating Public Sentiment in the Circular Economy through Topic Modelling and Hyperparameter Optimisation

### v1

@@ 3.3 Hyperparameter Optimisation for Topic Modelling
In this study, we selected hyperparameters to be optimised for three models according to the suggestions from original papers: BERTopic [27], CorEx [12], and LDA [42]. We have decided to collaboratively use BERTopic, CorEx, and LDA for topic modelling experiments in our research, based on their complementary analytical strengths, which together provide a robust methodological framework for dissecting complex changes in public sentiment. BERTopic leverages the powerful capabilities of large language model architectures, enabling it to handle vast amounts of unstructured text with a high degree of semantic understanding. This allows BERTopic to extract thematic clusters that reflect intricate details of public sentiment on the CE. CorEx employs an information-theoretic approach, capable of identifying subtle patterns and correlations within the data [12]. This makes the CorEx model particularly valuable in parsing the thematic structures behind public discourse, thus complementing the depth of BERTopic through meticulous detection of latent topics. LDA offers a probabilistic perspective on topic modelling, providing a broad overview of the topic distributions within the data. LDA’s strength lies in its ability to generalise across large texts, thus enabling a macro-level understanding of the popularity and shifts concerning CE topics. These three methods ensure a comprehensive exploration of topic distributions, capture the granularity and breadth of public discourse, and offer a multifaceted understanding of public sentiment towards CE on different online platforms.
@@
It is noted that random search and grid search are two popular approaches because they are simple and easy to implement. As shown in Table 1, our defined search spaces are not large compared to those in deep-learning research. Additionally, the three models have relatively lower computational complexities compared to large deep learning models. Therefore, we decided to use grid search to perform an exhaustive search to confirm the best hyperparameters, as using random search in such small search spaces does not make sense.
Moreover, it is worth drawing attention to the fact that topic modelling is a type of unsupervised learning. No single quantitative metric can comprehensively evaluate the results, pointing out the importance of human evaluation. Therefore, in our research, we will begin by comparing the hyperparameter optimisation results in Section 4.1 with human interventions. Subsequently, we will utilise the selected hyperparameter optimisation results to conduct further analysis (with visualisation) in the context of circular economy in Sections 4.2 $\sim$ 4.4.
@@
Table 1: The hyperparameters selected to be optimised and their corresponding search spaces.
@@ 4 Experiments and results
4.1 Hyperparameter Optimisation
The results of hyperparameter optimisation for three selected models (BERTopic, CorEx, and LDA) on three datasets (The Guardian, Reddit, and Twitter) are shown in Table 2. The coherence values (NPMI) presented in Table 2 indicate that BERTopic outperformed the other two models on two of three datasets. However, much of the research in topic modelling emphasises that evaluating the quality of topic modelling should not solely rely on quantitative metrics but also incorporate human evaluation. Therefore, we listed the topics obtained from three topic models with the best hyperparameters on three datasets in Sections 4.1.1 $\sim$ 4.1.3, respectively. In the following sections, we will compare, analyse, and discuss the topics obtained from different models, dataset by dataset, to decide which model will finally be selected for further analysis.
Model Dataset Best Parameters Best Coherence (c_npmi)
@@ 4.1 Hyperparameter Optimisation
Table 2: Best hyperparameter values and corresponding coherence scores for BERTopic, CorEx, and LDA models on three datasets.
@@
4.1.1 The Guardian
In Table 2, BERTopic achieved the best performance with a coherence value of 0.1669. To further investigate the performances of the three models, Tables 3 to 5 show the topics generated by three different models on the Guardian dataset. From our perspective, LDA topics are not desirable because the word ‘say’ appears and dominates in its three topics, but it is less meaningful for CE. Meanwhile, BERTopic and CorEx are relatively good; we can clearly see words such as ‘energy’, ‘gas’, and ‘power’, which are closely related to CE. Although we found ‘Australian’-related words in the topics, we manually checked our dataset and found that there are some news articles from the Australian news sector scraped by the API. Due to the large size of the dataset, we are not able to delete them manually, so we kept them and take the advantages of topic modelling to ignore them in the following analysis.
number value params_n_clusters params_n_components params_n_gram params_n_neighbors
@@ 4.1.1 The Guardian
Table 6: The top 5 hyperparameter sets for BERTopic on the Guardian dataset
@@
4.1.2 Reddit
@@
The BERTopic model (Table 8) identifies two topics. The first topic integrates core aspects of the circular economy such as ‘recycling’, ‘sustainable’ and ‘waste’ pointing towards operational and environmental concerns. The second topic seems to focus on the business and ideational perspectives, including terms like ‘business model’ and ‘circularity’, which suggest strategic to circular economy practices. BERTopic provides a better understanding of the contextual relationships between words compared to traditional LDA models because of the presence of the large language model BERT. By utilising BERT’s deep semantic understanding, it can link concepts such as environment and sustainability, generating related words such as ‘circular economy’ and ‘sustainable’. The word embedding of the BERT model allows the model to capture the link between specific activities such as ‘recycling’ and ‘plastic waste’ and the broader topic of circular economy.
@@
4.1.3 Twitter
While exploring user sentiments on the CE from Twitter, we utilised Twitter’s official API to collect posts containing the hashtag ‘Circular Economy’. However, starting in 2023, Twitter’s official API began to restrict the scope of user searches, allowing us only to access data from the past year. We conducted data cleansing on all collected data, retaining core information such as the content and timing of tweets. We also removed duplicates, replied and reposted tweets, ultimately preserving 3,922 English tweets for our model training.

### v2

@@ 3.3.2 MOO
In this research, we selected hyperparameters to be optimised for three models according to the suggestions from original papers: BERTopic [28], CorEx [12], and LDA [43]. We have decided to collaboratively use BERTopic, CorEx, and LDA for topic modelling experiments in our research, based on their complementary analytical strengths, which together provide a robust methodological framework for dissecting complex changes in public attention. BERTopic leverages the powerful capabilities of language model architectures, enabling it to handle vast amounts of unstructured text with a high degree of semantic understanding. This allows BERTopic to extract thematic clusters that reflect intricate details of public attention on the CE. CorEx employs an information-theoretic approach, capable of identifying subtle patterns and correlations within the data [12]. This makes the CorEx model particularly valuable in parsing the thematic structures behind public discourse, thus complementing the depth of BERTopic through meticulous detection of latent topics. LDA offers a probabilistic perspective on topic modelling, providing a broad overview of the topic distributions within the data. LDA’s strength lies in its ability to generalise across large texts, thus enabling a macro-level understanding of the popularity and shifts concerning CE topics. These three methods ensure a comprehensive exploration of topic distributions, capture the granularity and breadth of public discourse, and offer a multifaceted understanding of public attention towards CE on different online platforms.
@@
It is noted that random search and grid search are two popular approaches because they are simple and easy to implement. Table 1 outlines the search spaces we used, which were determined and suggested by relevant literature. These search spaces are relatively small compared to those typically encountered in deep learning research. Additionally, the three models employed in our work have lower computational complexity compared to large deep learning models. As a result, we opted for a grid search to conduct an exhaustive search and identify the optimal hyperparameters. One drawback of random search, in contrast to grid search, is the uncertainty introduced by its inherent randomness.
Moreover, it is worth drawing attention again to the fact that topic modelling is a type of unsupervised learning. No single quantitative metric can comprehensively evaluate the results, pointing out the importance of human evaluation and multi-objective optimisation. Therefore, in our research, we will begin by comparing the results in Section 4.1 with human interventions. Subsequently, we will utilise the selected hyperparameter optimisation results to conduct further analysis (with visualisation) in the context of CE in Sections 4.3.1 $\sim$ 4.5.1.
Table 1: The hyperparameters selected to be optimised and their corresponding search spaces for SOO.
@@ 4 Experiments and Results
4.1 Best Hyperparameter Values from SOO
The results of SOO for three selected models (BERTopic, CorEx, and LDA) on three datasets (The Guardian, Reddit, and Twitter) are shown in Table 2. The coherence values ($C_{\text{NPMI}}$) presented in Table 2 indicate that BERTopic outperformed the other two models on two of three datasets. Subsequently, we conducted a qualitative analysis of the topics produced by each model. This involved examining the top representative words for each topic to evaluate their interpretability and relevance. The goal was to observe whether the models, when optimally tuned, could generate meaningful and distinct topics that align with the semantic structures present in the data.
Table 2: Best Hyperparameter Values and Corresponding Coherence Scores for BERTopic, CorEx, and LDA on Three Datasets from SOO. $\uparrow$ means higher is better, $\downarrow$ means lower is better. Bold values represent the best performance for each dataset. The $C_{\text{NPMI}}$ metric ranges from $-1$ to $1$. A value closer to $1$ indicates higher topic coherence, meaning the words within topics are more semantically related. A value of 00 suggests no mutual information between words, while negative values indicate incoherence.
Model Dataset Best Parameters Best Coherence ($C_{\text{NPMI}}$) $\uparrow$
@@ 4.1 Best Hyperparameter Values from SOO
4.2 Best Hyperparameter Values from MOO
As discussed in Section 4.1, the SOO focused solely on maximising the coherence metric ($C_{\text{NPMI}}$) across three models. While this approach revealed that BERTopic consistently outperformed other models on two datasets, the researchers increasingly recognise the limitations of relying on a single metric for topic modelling evaluation. Trade-offs often exist between multiple objectives, such as coherence, topic diversity, and perplexity. To address this, we extended our optimisation to a MOO method to explore deeply.
The decision to apply MOO only to BERTopic was driven by several factors. In the summary, BERTopic has shown better performance under SOO, making it the most promising candidate for further optimisation. Unlike LDA and CorEx, BERTopic’s integration with the BERT language model allows for more complex semantic understanding, making it more adaptable to balancing multiple objectives simultaneously. From our perspective, BERTopic is designed to integrate with the language model, and is capable of handling texts from various scenarios, such as news articles and social media posts. Therefore, BERTopic provides a richer search space for MOO, allowing for fine-tuning across different dimensions.
On the other hand, CorEx and LDA did not demonstrate sufficient flexibility or performance gains in SOO to justify their inclusion in MOO as results showed. Compared to BERTopic, the other two approaches are less flexible. BERTopic can be considered a BERT-based topic modelling framework, allowing various computational components to be changed or their hyperparameters adjusted, providing greater flexibility to handle different tasks.
Table 3: Best Hyperparameter Values and Corresponding Scores for BERTopic on MOO (with Ideal Point-Based Process). $\uparrow$ means higher is better, $\downarrow$ means lower is better. Bold values represent the best performance for each dataset. Diversity ranges from 00 to $1$. Values closer to $1$ signify greater diversity among topics, indicating that topics are more distinct from one another. Perplexity ranges from $1$ to positive infinity. Lower perplexity values indicate better model performance, with values closer to $1$ being optimal.
Model Dataset Best Parameters $C_{\text{NPMI}}$uparrow Diversity $\uparrow$ Perplexity $\downarrow$
BERTopic The Guardian
n_gram=1,
n_clusters=20,
n_components=11,
n_neighbors=19
0.1381 0.9886 1.4469
Reddit
n_gram=1,
n_clusters=14,
n_components=13,
n_neighbors=15
-0.2627 0.8558 8.1827
Twitter
n_gram=1,
n_clusters=8,
n_components=7,
n_neighbors=15
-0.0111 0.9486 74.9202
By conducting MOO, we were able to identify optimal hyperparameter sets that strike a balance between these competing objectives. Table 3 presents the results from the best-performing models on The Guardian, Reddit and Twitter datasets, with optimal values for $C_{\text{NPMI}}$, diversity, and perplexity.
For the The Guardian dataset, the model reached a high $C_{\text{NPMI}}$ score of 0.1381, paired with a notable diversity score of 0.9886 and a low perplexity of 1.4469, indicating a well-rounded performance in formal content. In contrast, the Reddit dataset showed a lower $C_{\text{NPMI}}$ of -0.2627, though the model maintained strong diversity (0.8558) and handled informal text reasonably well, with a perplexity of 8.1827. On the Twitter dataset, BERTopic achieved a near-neutral $C_{\text{NPMI}}$ (-0.0111) with high diversity (0.9486), but a perplexity of 74.9202 reflects the challenges posed by short, context-specific tweets. Overall, BERTopic consistently performs well across multiple datasets, particularly in capturing diverse topics, even where coherence is more difficult to maintain.
4.3 The Guardian
4.3.1 SOO
@@
In Table 2, BERTopic achieved the best performance with a coherence value of 0.1669. To further investigate the performances of the three models, Table 4 to Table 6 show the topics generated by three different models on the Guardian dataset. From our perspective, LDA topics are not desirable because the word ‘say’ appears and dominates in its three topics, but it is less meaningful for CE. Meanwhile, BERTopic and CorEx are relatively good; we can clearly see words such as ‘energy’, ‘gas’, and ‘power’, which are closely related to CE. Upon analysing the topics, we observed that several Australian-related terms appeared prominently. This is due to The Guardian’s inclusion of articles from its Australian news section via the API. Recognising that Australia has a high level of engagement with the circular economy, we retained these articles to enrich our analysis. This inclusion aligns with our objective of capturing a global perspective on public attention towards the CE.
Table 7: The top 5 hyperparameter sets for BERTopic on the Guardian dataset.
number $C_{\text{NPMI}}$ params_n_clusters params_n_components params_n_gram params_n_neighbors
@@
Visualisation and Analysis
The remaining five topics each feature one or two words with significantly higher scores. For instance, in Topic 14, the score for ‘ev’ is notably higher than that of the other words. This is indicative of the strength of BERTopic as a language model that utilises contextual analysis to effectively summarise and comprehend abbreviations like ‘ev’. Other related terms, such as brands including ‘Tesla’ and ‘Nissan’, aid the model in more swiftly grasping the significance of ‘ev’. The descending order of scores also illustrates the prioritisation and segmentation of these words within the topic.
We acknowledge that BERTopic performs admirably in interpreting text within The Guardian dataset, aggregating relevant vocabulary under a unified topic. However, an analysis of the highest-scoring words from Topics 4 and 6 reveals that words of different grammatical forms often convey essentially the same meaning. Both ‘recycling’ and ‘recycle’ in the text reflect the concept of resource recovery, while ‘sustainability’ and ‘sustainable’ repeatedly express the notion of environmental sustainability. We believe that by leveraging the strengths of language models, including part-of-speech normalisation and selective word comparison, BERTopic can offer researchers a more comprehensive vocabulary pertaining to each topic.
The frequencies and corresponding times of the seven topics are presented in Figure 6. Topic 1 shows a growing trajectory, rising significantly starting around 2017. This upgrade is timed to coincide with key international developments such as the 2016 Paris Agreement, highlighting the growing focus in the UK news on the synergies between renewable energy adoption and carbon emissions. Topic 15 experienced a surge in frequency around 2019. The topic continues to gain visibility, and this rise may be related to increased public concern about chemical pollution incidents and environmental health. The increase in frequencies of all topics post-2015 suggests an overall growing public and journalistic interest in environmental issues. This period seems to be pivotal in shaping public attention to sustainability and environmental topics.
Figure 7 presents a heatmap of the similarity matrix for topics generated by the BERTopic model on The Guardian dataset. These seven topics correspond to the ones shown in Figure 5. Deeper shades represent higher cosine similarity, whereas lighter shades indicate lower similarity. Since we get more topics in the Guardian dataset than in the Reddit and Twitter datasets, we believe the heatmap in Figure 7 facilitates a comparative analysis of the interrelationships among the topics.
According to this scatter plot, we can notice that the cluster of Topic 14 (pink colour) is somewhat separated from the rest of the topic clusters. This means that the news around electric vehicles in The Guardian is more of a dedicated discussion to this area, focusing on key industry players such as Tesla. While this topic may not be at the centre of most discussions in The Guardian news, it can be a good representation of the impact of innovation trends, and CE being carried out in transport. Additionally, although Topic 15 (grey colour) is similar to other clusters, it has a smaller size which may reflect the specificity of discussions around toxic chemicals. The size of clusters in this topic shows there is a specialised and smaller number of news about chemical terms and EPA in The Guardian dataset.
In the field of topic modelling, word clouds serve as a visual heuristic method, displaying terms that occupy central nodes in the discourse network [68]. Figure 9 shows the topic vocabulary graph cloud generated by BERTopic on The Guardian dataset. The visual prominence of terms such as ‘sustainability’, ‘sustainable’, ‘recycling’, and ‘energy’ within the cloud is indicative of their pivotal role in the corpus, highlighting the dataset’s focus on environmental conservation, renewable energy, and the principles of the CE. These terms underscoring the centrality of sustainable energy in CE are captured within The Guardian news. Surrounding these core terms are secondary and substantial, concepts such as ‘ev’, ‘electric’, and ‘vehicle’, which given their relative visual weight (size), suggest substantive topics that concern technological advancements in transportation and the transition. Moreover, the adjacency of concepts such as ‘ethical’, ‘consumer’, ‘growth’, and ‘future’ alongside ’sustainability’ within the visualisation elucidates an interplay of notions.
4.3.2 MOO
The topics generated by BERTopic on The Guardian dataset, optimised via MOO, provide a more nuanced view of the various discussions related to the CE and other significant topics because the multi-objective optimisation balances coherence, diversity, and perplexity. This balance allows the model to capture a broader spectrum of topics, uncovering subtle and diverse topics that might be overlooked when focusing on a single objective. By optimising multiple objectives simultaneously, we enable the discovery of richer and more comprehensive topic representations within the data. Notably, Topic 2, which includes terms such as ‘renewable’, ‘energy’ and ‘uk’ focuses heavily on UK environmental and energy-related issues. This aligns well with the core concerns of the CE, showcasing how MOO helps capture a broader and more diverse range of topics compared to single-objective optimisation. The inclusion of relevant terms such as ‘solar’ and ‘tory’ suggests an intersection of political and environmental discussions, providing a richer context for researchers examining the policy implications of climate change.
Another critical topic, Topic 4, brings together terms like ‘sustainability’, ‘resource’ and ‘future’ directly relating to the CE. This topic reflects broader discussions about sustainable development goals and the future of resource management, emphasising the model’s ability to capture key aspects of CE debates. By incorporating terms such as ‘deforestation’ and ‘energy’, MOO allows the model to go beyond narrow environmental concerns, offering insights into the strategic planning and forward-looking goals of CE.
The diversity of topics presented in Table 9 also reveals insights into current social and political affairs, such as Topics 1 and 3, which focus on UK politics, and Topic 11, which highlights the impact of the COVID-19 pandemic. These topics provide a wider socio-political context, reflecting the broader scope of media coverage in The Guardian. Overall, MOO enables a more comprehensive extraction of topics, providing researchers with a broader spectrum of relevant topics related to the CE and beyond, ensuring that key dimensions of current societal challenges are well-represented in the analysis.
On The Guardian platform, the results of SOO and MOO highlight different optimisation focuses. SOO primarily maximises topic coherence, focusing on the semantic consistency of the generated topics. Topic models optimised by SOO, such as BERTopic, show relevance in the generated topic words, with terms like ‘energy’, ‘gas’, and ‘emission’ closely tied to environmental sustainability. However, due to its sole focus on coherence, the number of generated topics is small (e.g., 2 topics), resulting in limited diversity.
In contrast, MOO generates a more diverse set of topics by optimising coherence, diversity, and perplexity simultaneously. The BERTopic model in MOO produces 20 topics, covering a wider range of circular economy topics, such as ‘recycling’, ‘renewable energy’, and ‘sustainability’. This demonstrates that the MOO method maintains a high level of semantic coherence while enhancing topic diversity. Moreover, the lower perplexity in MOO indicates that the model is more stable when predicting unseen text. Overall, the MOO model increases topic diversity and model robustness while maintaining high coherence, making it suitable for analysing complex text data.
Table 9: The topic terms obtained from the BERTopic model with the best hyperparameter values on The Guardian dataset (MOO).
Topic Topic Term
0 australia, australian, nsw, bst, queensland, sydney, gmt, labor, election, people
1 corbyn, tory, mp, uk, brexit, conservative, eu, election, britain, question
2 renewable, energy, tory, uk, britain, solar, consumer, carbon, conservative, emission
3 cop27, cop26, carbon, emission, world, global, energy, international, 20230, renewable
4 sustainability, sustainable, resource, future, create, ethical, develop, challenge, organisation, environment
5 rio, world, sdgs, earth, resource, environment, international, global, deforestation
6 meat, food, farming, feed, farmer, beef, produce, dairy, crop, agriculture, farm
7 car, emission, vehicle, petrol, drive, electric, charge, diesel, driver, carbon
8 recycling, recycle, recyclable, rubbish, waste, plastic, landfill, reuse, circular, litter
9 feel, moment, need, want, share, use, think, tell, say, medium
10 candidate, trump, biden, voter, democrat, election, official, senator, democrats, republicans
11 covid, coronavirus, vaccine, vaccination, lockdown, world, gmt, outbreak, pandemic, die
12 ftse, inflation, eurozone, fed, uk, gmt, market, investor, gdp, trading
13 arsenal, player, sport, liverpool, club, league, final, chelsea, goal, team
14 housing, property, house, landlord, building, site, tenant, land, build, resident
15 clothe, wear, garment, fashion, clothing, textile, retailer, consumer, shirt, fabric
16 epa, pfas, chemical, toxic, pesticide, environmental, federal, emission, pollution, lead
17 ukraine, ukrainian, russia, putin, kyiv, kremlin, eu, world, russian, attack
18 golf, tournament, final, putt, forehand, mcilroy, tee, open, player, backhand
4.4 Reddit
4.4.1 SOO
@@
The BERTopic model (Table 10) identifies two topics. The first topic integrates core aspects of the circular economy such as ‘recycling’, ‘sustainable’ and ‘waste’ pointing towards operational and environmental concerns. The second topic seems to focus on the business and ideational perspectives, including terms like ‘business model’ and ‘circularity’, which suggest strategic to circular economy practices. BERTopic provides a better understanding of the contextual relationships between words compared to traditional LDA models because of the presence of the language model BERT. By utilising BERT’s deep semantic understanding, it can link concepts such as environment and sustainability, generating related words such as ‘circular economy’ and ‘sustainable’. The word embedding of the BERT model allows the model to capture the link between specific activities such as ‘recycling’ and ‘plastic waste’ and the broader topic of circular economy.
@@
Visualisation and Analysis
The word cloud (Figure 13) not only captures the direction of Reddit user’s discussions about CE but also provides insight into the collective focal points of the community, thereby reflecting the broader public engagement and attention towards sustainability.
4.4.2 MOO
As shown in the Table 13, topic 1 includes terms such as ‘circular’, ‘economy’ and ‘environment’ demonstrating the model’s ability to capture the broader intersection between environmental issues and economic strategies. This reflects the community-driven nature of Reddit, where discussions often highlight the socio-economic implications of environmental policies. The inclusion of words like ‘economic’ and ‘environment’ suggests a nuanced understanding of the relationship between sustainable practices and economic outcomes.
Topic 3 focuses on recycling, with terms such as ‘recycling’, ‘reuse’ and ‘reusable’. Topic 3 highlights the core components of the CE. This shows how MOO effectively captures the practical aspects of waste management and resource reuse, which are crucial topics in any CE discourse. By extending beyond simple coherence optimisation, MOO reveals the depth of user-generated content on Reddit, which often centres on community-led initiatives and grassroots movements.
The diversity of topics in Table 13 demonstrates how MOO can uncover not only the technical aspects of CE but also the societal and economic dimensions. For instance, Topic 2 brings together terms like ‘eu’, ‘circular’ and ‘sustainable’, indicating ongoing discussions about European sustainability and resource management policies. Overall, the MOO results provide a well-rounded analysis of CE-related discussions, enabling researchers to better understand the key issues being debated on platforms like Reddit.
On the Reddit platform, the topic model generated by SOO optimisation performs poorly in terms of coherence (negative C NPMI value), likely due to the scattered and informal nature of user discussions on Reddit. The SOO results show a small number of topics (e.g., 2), which may not fully capture the breadth of public discussion on the platform. Additionally, SOO’s topic words contain more redundant information, such as ‘say’ and ‘year’, reflecting SOO’s limitations when dealing with dispersed data.
MOO optimisation, on the other hand, generates more topics (14) and significantly improves the diversity score. This suggests that, despite the scattered nature of user discussions on Reddit, MOO can still capture a more diverse range of topic structures. Furthermore, although MOO’s consistency score is lower, the improvement in perplexity indicates that the model performs better when dealing with unstructured and varied content. Therefore, MOO is better at balancing consistency and diversity on Reddit, allowing it to capture more of the platform’s public concerns.
Table 13: The topic terms obtained from the BERTopic model with the best hyperparameter values on the Reddit dataset (MOO).
Topic Topic Term
0 sustainability, sustainable, lca, questionnaire, circular, survey, feedback, project, idea, question
1 circular, economy, environment, economic, sustainability, emission, future, energy, resource, innovation
2 eu, circular, european, sustainable, europe, 2030, ubi, environment, consumer, project
3 recycling, recycle, reuse, reusable, circular, sustainable, upcycled, idea, use, packaging
4 recycling, recycle, recyclate, plastic, microplastic, environmental, circular, packaging, sustainable, waste
5 podcast, circular, sustainable, sustainability, resource, circularity, consumption, lecture, opportunity, innovation
6 circular, clothing, outfit, clothe, garment, fashion, textile, survey, wear, feedback
7 recycling, environment, recycle, circular, energy, greenwashing, sustain, cyclical, environmentalist, ipcc
8 circular, circularity, opendesign, modeindustrie, design, video, eco, materiali, biomimicry, animation
9 repair, repairable, fix, repairer, repairability, gadget, ifixit, device, electronic, check
10 recycle, recycling, recycled, recyclemore, renewable, zerowaste, circular, reuse, eco, waste
11 recycle, recycling, recyclable, ev, solar, pv, energy, electrocatalyst, battery, circular
12 webinar, circular, upcoming, register, coursera, hub, registration, event, workshop, course
13 macarthur, circular, ellen, toolkit, report, programme, foundation, interested, 2020, project
4.5 Twitter
4.5.1 SOO
While exploring user attention on the CE from Twitter, we utilised Twitter’s official API to collect posts containing the hashtag ‘Circular Economy’. However, starting in 2023, Twitter’s official API began restricting the scope of user searches, limiting us to accessing data from the past year. We conducted data cleaning on all collected data, retaining core information such as the content and timing of tweets. We also removed duplicates, replies, and reposted tweets, ultimately preserving 3,922 English tweets for our model learning.